PERMANENT ELECTRON ELECTRIC DIPOLE MOMENT SEARCH IN THE X^{3}Δ_{1} GROUND STATE OF TUNGSTEN CARBIDE MOLECULES

\underline{JEONGWON LEE}, JINHAI CHEN, and AARON LEANHARDT, Department of Physics, University of Michigan, Ann Arbor, MI 48109.

We are developing an experiment to search for the permanent electric dipole moment (EDM) of the electron using the valence electrons in the X^{3}Δ_{1} ground state of Tungsten Carbide (WC) molecules. Currently, we are detecting the molecules by Laser Induced Fluorescence spectroscopy at ∼75cm downstream of a pulsed ablation beam source. We have a detection rate of ∼10 ^{182}W^{12}C molecules/second in X^{3}Δ_{1}, v”=0, J”=1 state with geometric detection efficiency of 0.004. A continuous WC molecular beam is under development. Additionally, preliminary measurements of the ^{183}W^{12}C hyperfine structure will be presented.